OpenStudy (lcnettech):

will give medal .ellas two puppys weigh a total of 37 oz.one puppy weighs 3 oz more than the other.how much does the lighter puppy weigh.use the guess and test strategy to answer.please help.

OpenStudy (lcnettech):

can someone please help?

OpenStudy (texaschic101):

guess and test strategy ?? puppy A + puppy B = 37 A + B = 37 A = B + 3 (B + 3) + B = 37 2B + 3 = 37 2B = 37 - 3 2B = 34 B = 34/2 B = 17 A + B = 37 A + 17 = 37 A = 37 - 17 A = 20 check.. puppy A + puppy B = 37 20 + 17 = 37 37 = 37 (correct) The lighter puppy (B) weighs 17 oz. Thats the answer...but I think for the guess and test strategy...if you guess 1 puppy was 20 and 1 was 23.....20 + 23 = 43..that would be too high. If you guess 1 was 15 and 1 was 18....15 + 18 = 33....too low. And I think you can just sub in guesses until you find the correct one which is both puppies weighing 20 and 17 oz, with the lighter one being 17 oz.

OpenStudy (lcnettech):

ok thanks. :)